A shell acquires the initial velocity v = 320 m/s, having made n = 2.0 turns inside the barrel whose length is equal to l = 2.0 m. Assuming that the shell moves inside the barrel with a uniform acceleration, find the angular velocity of its axial rotation at the moment when the shell escapes the barrel.
1
Expert's answer
2020-10-13T09:36:05-0400

"\left.\begin{matrix}
 {\omega =2 \pi \nu} \\ 
 \left.\begin{matrix}
 \nu =\frac{n}{t} \\ 
 t=\frac{l}{v} 
\end{matrix}\right\}
\Rarr
\nu =\frac{nv}{l} 
\end{matrix}\right\}
\Rarr
\omega =\frac{2 \pi nv }{l}"

"\omega =\frac{2\cdot 3.14\cdot 2\cdot320 }{2}\approx2000" ("\frac{rad}{s}" )